### v1

@@ Abstract
This paper extends the Isotonic Mechanism from the single-owner to multi-owner settings, in an effort to make it applicable to peer review where a paper often has multiple authors. Our approach starts by partitioning all submissions of a machine learning conference into disjoint blocks, each of which shares a common set of co-authors. We then employ the Isotonic Mechanism to elicit a ranking of the submissions from each author and to produce adjusted review scores that align with both the reported ranking and the original review scores. The generalized mechanism uses a weighted average of the adjusted scores on each block. We show that, under certain conditions, truth-telling by all authors is a Nash equilibrium for any valid partition of the overlapping ownership sets. However, we demonstrate that while the mechanism’s performance in terms of estimation accuracy depends on the partition structure, optimizing this structure is computationally intractable in general. We develop a nearly linear-time greedy algorithm that provably finds a performant partition with appealing robust approximation guarantees. Extensive experiments on both synthetic data and real-world conference review data demonstrate the effectiveness of this generalized Isotonic Mechanism.
@@ 1 Introduction
In recent years, major machine learning conferences such as NeurIPS and ICML have faced a concerning decline in the quality of peer review, which poses a significant challenge to the global machine learning community [16, 5, 24]. In particular, the NeurIPS 2014 experiment showed that $49.5\%$ of the papers accepted by one committee would be rejected by another [17, 8, 16]. This inconsistency probability was $50.6\%$ for NeurIPS 2021 [6]. This worrying trend is, in part, due to the surge in submission volumes [20]. For example, NeurIPS 2022 witnessed a record-breaking 10,411 submissions. In contrast, as demonstrated in [15, 19, 21], the growth rate of qualified reviewers lags significantly behind this surge in submissions.
Figure 1: An example of an author-paper ownership set shown as a bipartite graph. An edge between an individual and a paper means that this individual is the author of the paper.
To mitigate this issue, there has been a progressive trend to propose various strategies aimed at enhancing the peer review process in machine learning (see, e.g., [20]). These strategies include improved reviewer assignments, the incorporation of additional questions, and many others [9, 14, 25, 13, 12]. Among these, an emerging approach employs mechanism design techniques to solicit private information from authors, thereby enabling a more accurate estimation of review scores [22, 23]. This approach, termed the “Isotonic Mechanism”, solicits authors of multiple submissions to rank their papers based on their perceptions. The adjusted review scores maintain consistency with the author-provided ranking while remaining as close as possible to the original review scores.
To formally describe this method, imagine that an author, say, Alice, submits $n$ papers to a conference. Right after submission, Alice under the Isotonic Mechanism provides a ranking $\pi$ in the form of a permutation of $1,2,\ldots,n$ that sorts her papers in descending order of quality. Letting $y_{1},\ldots,y_{n}$ denote the (average) review scores of the $n$ papers, the Isotonic Mechanism yields adjusted review scores that are the solution to the following convex optimization program:
@@
Under certain utility assumptions, [22] shows that Alice is incentivized to provide the true ranking under the Isotonic Mechanism, which in turn improves the estimation accuracy of the “ground-truth” review scores. The improvement is substantial when the number of submissions is large and the review scores contain a high level of noise. Moreover, [23] proves that the ranking of the submissions is the most fine-grained information that can be truthfully elicited from the author.
The Isotonic Mechanism was recently experimented at ICML 2023111See https://icml.cc/Conferences/2023/CallForPapers.. However, this method in its current form cannot fully adapt to practical settings wherein papers have multiple authors. Specifically, it is not clear how the rankings of Alice’s co-authors would affect the utility of Alice. A more challenging situation is where different papers partially overlap in authorship, as illustrated in Figure 1. In such a case, Alice’s utility might be affected by other authors’ rankings that include papers Alice did not contribute, if one applies the Isotonic Mechanism to every author’s submissions and subsequently averages the adjusted review scores for each paper. This is an unintended effect that we aim to eliminate in an ideal mechanism.
In this paper, we introduce a generalized Isotonic Mechanism to address the challenge of overlapping authorship in large machine learning conferences. Our approach begins by partitioning the author-paper ownership sets into blocks of papers, ensuring that each block shares a common set of authors. We then apply the Isotonic Mechanism to each block, soliciting paper rankings from each common author, and outputting a weighted average of all adjusted review scores from all calls of the Isotonic Mechanism for this block. Under certain conditions, this generalized mechanism is marked by the following key characteristic:
@@
Truthfully reporting rankings by all authors forms a Nash equilibrium under the generalized Isotonic Mechanism.
Along with other results, the truthfulness result is formally presented in Section 3. This result suggests that the optimal strategy for an author to maximize her utility is to report the ground-truth ranking, given that all the other authors do likewise. Its proof involves a novel technique concerning majorization. Additionally, we demonstrate that this Nash equilibrium is the most favored, as any other potential Nash equilibrium would result in lower utility for every author.
As is conceivable, the statistical performance of this generalized Isotonic Mechanism depends on the partition. For example, a partition for Figure 1 can be simply $\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$, which share a single common author $\{a_{3}\}$, or can be $\{p_{1},p_{2},p_{3}\},\{p_{4}\}$, and $\{p_{5},p_{6}\}$, which share $\{a_{1},a_{2},a_{3}\},\{a_{3}\}$, and $\{a_{3},a_{4}\}$, respectively. To find high-quality author-paper ownership partitions, we define a natural class of criteria to assess the quality of a partition. Within this class of criteria, we obtain our second main result concerning the generalized Isotonic Mechanism:
@@
A simple linear-time greedy algorithm can obtain a partition of the author-paper ownership set for the generalized Isotonic Mechanism that is simultaneously near-optimal for every criterion in the class.
This result is formally stated in Section 4. In the same section, we show that it is NP-hard to find the exact optimal partition even for a certain simple criterion. In contrast, the greedy algorithm is computationally efficient while achieving constant-ratio approximation simultaneously for most natural criteria.
We validate this mechanism through a series of experiments on both ICLR review data and synthetic data. Our experiments exhibit that this generalized Isotonic Mechanism enhances estimation performance relative to the original review scores, a benefit observed consistently across various experimental settings.
2 Notation and Preliminaries
This paper extends the problem setup in [23] to a general model of overlapping ownership. There are $m$ owners and $n$ items. The $j$-th owner owns a set of items $\mathcal{I}^{j}\subseteq[n]=\{1,2,\dots,n\}$ and let $|\mathcal{I}^{j}|=n_{j}$. The ownership relation can be visualized as a bipartite graph (e.g., see Figure 1 and their formal connections in Section 6.2). For each item $i=1,\ldots,n$, let $R_{i}$ be its ground-truth quality and
@@ 2 Notation and Preliminaries
be a noisy rating of its quality by reviewers. As a friendly reminder of our notation, $i$ and $\mathcal{I}$ are used to denote item index and item set, respectively. Denote by $\bm{z}=(z_{1},z_{2},\dots,z_{n}),\bm{R}=(R_{1},R_{2},\dots,R_{n})$ and $\bm{y}=(y_{1},y_{2},\dots,y_{n})$, respectively, the vector of noise, ground-truth item scores and review scores. Let $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ be the $j$-th owner’s utility function on the set of items she owns. We adopt the same assumptions (listed below) as [23] about owner knowledge, noise distribution and utility.
Assumption 2.1 (Informed Owners [23]).
For each $j$, the $j$-th owner has sufficient knowledge of the ground-truth item quality $\bm{R}$ to determine the true ranking of her own items $\mathcal{I}^{j}$.
Assumption 2.2 (Exchangeable Noise Distribution [23]).
@@
The review noise vector $\bm{z}=(z_{1},\dots,z_{n})$ follows an exchangeable distribution in the sense that $(z_{1},\dots,z_{n})$ has the same probability distribution in $\mathbb{R}^{n}$ as $\rho\circ\bm{z}:=(z_{\rho(1)},\dots,z_{\rho(n)})$ for any permutation $\rho$ of $1,2,\dots,n$.
Assumption 2.3 (Convex Utility [23]).
@@
For simplicity of presentation, we focus on the most standard knowledge partition specified by the $n!$ permutations of the ordinal order of the item qualities. For Assumption 2.3, the utility function can be slightly relaxed to a non-separable utility function $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ that satisfies Schur-convexity, namely,
$(x_{i}-x_{i^{\prime}})\left(\frac{\partial\mathrm{U}^{j}(\bm{x})}{\partial x_{i}}-\frac{\partial\mathrm{U}^{j}(\bm{x})}{\partial x_{i^{\prime}}}\right)\geq 0,$
for all $i,i^{\prime}\in[n_{j}]$. Notably, Assumption 2.3 does not require ${U}^{j}={U}^{j^{\prime}},\forall j\neq j^{\prime}$, so each owner may have different utility on the same scoring of their items. We refer the readers to [23] for the detailed discussions and justification for these assumptions.
At a high level, our goal is to design a scoring mechanism $\mathcal{M}$ to estimate item qualities $\bm{R}$. The mechanism is expected to mitigate noises in review scores $\bm{y}$ by incentivizing as many owners to truthfully share as much information on their item qualities as possible. The key to this special kind of mechanism is the design problem of a statistical estimator $\widehat{\mathrm{R}}_{\mathcal{M}}$ about the item qualities $\bm{R}$. To improve on the observed review score $\bm{y}$, the estimator will additionally account for the ranking information elicited from the owners $\{\pi^{j}\}_{j=1}^{m}$. Motivated by the practical applications of conference peer review, we require the mechanism to proceed in the following stages:
  1. 1.
    Before Submission: Each owner is notified of the scoring mechanism and obtains some knowledge on the qualities of the items that she owns.
  2. 2.
    During Submission: Each owner $j$ reports a ranking $\pi^{j}$ of her items $\mathcal{I}^{j}$ [^2] If an owner does not report a ranking, a uniformly random ranking will be used. As will be shown later, this design ensures rational owners participate in the mechanism. It is also without loss of generality to assume that owners report the full ranking of their items, and we defer the discussion to Remark 1.2.
  3. 3.
    After Submission: Once the review score $\bm{y}$ is realized, the mechanism accordingly estimates the item qualities as $\widehat{\bm{R}}=\widehat{\mathrm{R}}_{\mathcal{M}}(\{\pi^{j}\}_{j=1}^{m};\bm{y})$. Each owner $j$ then derives utility $\mathrm{U}^{j}\left([\widehat{R}_{i}]_{i\in\mathcal{I}^{j}}\right)$.
We use the standard Bayesian game framework [10] to model the owners’ strategic decision-making under the uncertainty of realized review scores $\bm{y}$. We say a profile of owners’ report $\{\pi^{j}\}_{j=1}^{m}$ forms a Bayes-Nash Equilibrium (NE) under mechanism $\mathcal{M}$  [^3] It suffices to work with this special set of pure-strategy NEs in most part of this paper and we defer the additional notations for the more general definition of mixed-strategy NEs to Appendix 1. if for any owner $j\in[m]$, given others’ report $\pi^{-j}=\{\pi^{j^{\prime}}\}_{j^{\prime}\neq j}$, $j$’s expected utility by reporting $\pi^{j}$ is no worse than reporting any other possible ranking $\widetilde{\pi}^{j}$ or, mathematically,
[…]
We say a mechanism $\mathcal{M}$ is truthful if every owner $j$ reporting the true ranking $\pi^{j}$ of her items forms a Bayes-Nash equilibrium under scoring mechanism $\mathcal{M}$.
3 Truthfulness under Completely Overlapping Ownership
The Isotonic Mechanism [23] in case of a single author uses any reported ranking $\pi$ from the owner and the realized review scores $\bm{y}$ for his items to solve for the item quality estimates,

### v3

@@ Abstract
Motivated by the problem of improving peer review at large scientific conferences, this paper studies how to elicit self-evaluations to improve review scores in a natural many-to-many owner-item (e.g., author-paper) situation with overlapping ownership. We design a simple, efficient and truthful mechanism to elicit self-evaluations from item owners that can be used to calibrate their noisy review scores in the existing evaluation process (e.g., papers’ review scores from peers).
Our approach starts by partitioning the owner-item relation structure into disjoint blocks, each sharing a common set of co-owners. We then elicit the ranking of items from each owner and employ isotonic regression to produce adjusted item scores, aligning with both the reported rankings and raw item review scores. We prove that truth-telling by all owners is a payoff dominant Nash equilibrium for any valid partition of the overlapping ownership sets under natural conditions. Moreover, the truthfulness depends on eliciting rankings independently within each block, making block partition optimization crucial for improving statistical efficiency. Despite being computationally intractable in general, we develop a nearly linear-time greedy algorithm that provably finds a performant block partition with appealing robust approximation guarantees. Extensive experiments on both synthetic data and real-world conference review data demonstrate the effectiveness of our mechanism in a pressing real-world problem.
@@ 1 Introduction
In recent years, major AI and machine learning conferences such as NeurIPS, ICML, and ICLR have faced a concerning decline in the quality of peer review, posing a significant challenge to the global machine learning community [21, 6, 39, 24]. In particular, the NeurIPS 2014 experiment showed that $49.5\%$ of the papers accepted by one committee would be rejected by another [22, 11, 21]. This inconsistency probability was $50.6\%$ for NeurIPS 2021 [9]. This troubling trend can be largely attributed to a structural imbalance between the surge in submission volumes and lagged growth of the number of qualified reviewers [32]. As demonstrated by Figure 1, many of these conferences have handled around 10,000 full paper submissions recently, reflecting an almost ten-fold growth over the past decade, whereas a large portion of the reviewers are inexperienced researchers (e.g., junior graduate students) [33, 34, 31]. Similar trends are seen across various research communities [27, 20, 14, 8], highlighting a systemic issue in the today’s peer review systems. To mitigate this issue, there have been progressive research efforts to improve peer review processes. Some approaches employ machine learning and optimization techniques to improve reviewer assignments, reduce bias, and automate review procedures, while others adopt an economic approach modeling the incentives of participants to encourage high-quality reviews (see, e.g., a survey by [32]).
Figure 1: The number of paper submissions in major machine learning and related conferences from 2014 to 2024.
Motivated by pressing concerns as mentioned above, this paper proposes a mechanism that blends together statistical and economic design, and can effectively incentivize authors to share evaluations of their own papers to calibrate peer review scores. Conceptually, the motivation is natural: if the reviewer pool cannot supply sufficient information for all the submissions, conference organizers could resort to other sources — in our case, the authors themselves. However, while self-criticism is a merit in science, self-evaluation has been largely overlooked due to the obvious conflicts of interest. That is, authors may be reluctant to provide honest assessments, if revealing a mediocre rating of their own work puts them at disadvantage. Consequently, estimations from manipulated data may be hardly useful. This presents a fundamental design challenge:
How can statistical estimation processes effectively elicit data from strategic actors?
In particular, is it possible to balance between statistical efficiency and incentive compatibility?
Examining the above question within a clean and stylized model of overlapping ownership, this paper embarks on a path to bring the mechanism design perspective to statistical estimation and proposes an approach, termed the Isotonic Mechanism, that demonstrates a definitive solution to this design challenge. In particular, while most of the learning algorithm designs have focused on improving the statistical efficiency, this work takes one step forward to further trace back the source of data. Specifically, through a careful balance of statistical efficiency and incentive compatibility, we design a mechanism that enables the elicitation and use of information traditionally deemed inaccessible. We view the demonstration of this feasibility in a high-stakes application as an important conceptual contribution, which we hope to spark a broader research agenda at the interplay of statistical and incentive design. Furthermore, given the prevalence of many-to-many owner-item relationships with overlapping ownership structures in real-world settings, we anticipate that our framework can be similarly applied to refine evaluation and award selection across a variety of domains — including internal team projects within large organizations, collaboratively produced artworks, and properties co-listed by multiple real estate agents.
To elicit an owner’s evaluation about her owned items, a key challenge to begin with, even putting incentive issues aside, is to determine what kind of data could be elicited from owners with reasonable accuracy. One approach might be to ask for highly fine-grained data, such as the owners’ own evaluation values (called item scores henceforth) for their items. However, even if owners are completely honest, they may lack the precise knowledge needed to provide reliable scores. Alternatively, one could request more general information (e.g., asking owners to identify their favorite item), but this may not offer sufficient value for improving review scores. The challenge lies in finding an effective middle ground—eliciting data that is both accurate and useful. To this end, we propose focusing on the ranking of an owner’s items (e.g., an author’s papers). Such relative information tends to be less noisy than absolute measures and has been successfully applied in various fields, such as learning from human preferences [44, 3].
Next comes our core scientific question: is it possible to truthfully elicit ranking data from owners to improve the statistical efficiency of any pre-given noisy review scores (e.g., papers’ original peer review scores)? Our starting point is an earlier work [36] which showcased the possibility in a much simplified situation with only a single owner, say, Alice. For this stylized case, Su [36] considered a natural statistical model for review score estimation and proved that truthful ranking elicitation is possible under natural owner utility assumptions. To formally describe this method, suppose Alice is the owner of $n$ items (e.g., a single author for $n$ papers). Letting $y_{1},\ldots,y_{n}$ denote any pre-given noisy review scores of the $n$ items (e.g., average reviewer scores for Alice’s single-authored $n$ papers). The mechanism elicits from Alice a ranking $\pi$ in the form of a permutation of $1,2,\cdots,n$ that sorts her items in descending order of quality, and then use this elicited ranking to compute adjusted review scores that are the solution to the following convex optimization program:
@@
This optimization yields the well-known isotonic regression [5], and its solution ensures consistency with the owner-provided ranking while remaining as close as possible to the original review scores in the least squares sense. Under natural statistical modeling of the pre-given review scores, this approach provably improves items’ score estimation. Moreover, it also guarantees truthful elicitation of Alice’s ranking, under standard utility and prior knowledge assumptions. Notably, the latter is a non-trivial property. Specifically, it crucially hinges on geometric properties of isotonic regressions, and would not hold for other estimation methods (e.g., switching from $l_{2}$-norm above to $l_{1}$-norm). This illustrates the importance of designing appropriate statistical estimation methods that can align with incentives.
Figure 2: An example of an owner-item ownership set shown as a bipartite graph. An edge between an owner and an item indicates that this individual owns the item.
This work seeks to address the more general situation where many individuals owners many items, with overlapping ownership (e.g., overlapping paper authorship). This strict generalization exhibits multiple new challenges. First, since Alice’s co-owners will also submit rankings for (some of) Alice’s items, Alice would have to account for the effect of her co-owners’ data on the utility of Alice. This intricacy is due to the fundamental difference between single-agent and multi-agent decision making setup, and it leads to our different solution concepts, generalizing from Bayesian optimal decision to Bayes Nash equilibrium. Second, an even more challenging issue is that different items can partially overlap in ownership, as illustrated in Figure 2. This is an especially common situation in today’s popular machine learning conferences. In such cases, Alice’s utility may be affected by other co-owners’ rankings on items Alice does not own. Naive adaptation of the previous Isotonic Mechanism designed for single-owner situations can lead to serious incentive issues. For instance, one natural approach is to apply the Isotonic Mechanism for every owner’s items and output final item scores by averaging their adjusted review scores. Unfortunately, it is easy to find undesirable examples with untruthful behaviors under this mechanism (see Section 3.2) because some co-owners would misreport their items’ ranking information in order to improve their own utilities. Therefore, a more principled “truth serum” is needed to effectively incorporate self-evaluations from a complex network of overlapping ownership into the reviewing mechanism.
To address the aforementioned challenges, this paper develops an approach that advances the earlier design using novel techniques from both algorithm design and mechanism design. Specifically, the mechanism first resorts to an algorithmic preprocessing procedure that partitions the owner-item ownership structure into blocks of items such that each block shares a common set of owners. It then uses the ranking information elicited from the common owners in each block to calibrate the pre-given review scores of their items. Under certain conditions, this proposed mechanism is marked by the following key characteristic:
@@
Truthfully reporting rankings by all owners forms a Nash equilibrium under the proposed mechanism.
Along with other results, the main result of truthfulness property is formally presented in Section 3.111We make a remark on practical implementation of our mechanism. To guarantee truthfulness, it suffices for the designer to commit to only use the partial ranking information within the partitioned blocks. Thus, to ease the practical implementation, the designer can elicit every owner’s full ranking of all her items but commit to use only part of the ranking information for the mechanism. In paper review applications, the mechanism designer is typically the organizing committee of large conferences. They do have such power of commitment, since their decisions are usually audited by steering committees as well. This result significantly extends the previous truthfulness guarantee of the earlier work [36] to cases with overlapping ownership and suggests that the optimal strategy for each owner to maximize her utility is to report the ground-truth ranking, given that all the other owners do likewise. In particular, our proof of this new result involves a novel technique concerning majorization in the ordering of item scores with respect to both the pre-given reviews’ noise and other owners’ reports. Moreover, we show the mechanism encourages truthfulness behaviors in several ways. From a game-theoretical perspective, this Nash equilibrium of truthfulness is the most favored, as any other potential Nash equilibrium results in no better utility for any of the owners. From empirical studies, the mechanism induces a common-interest game and the human subjects are found to always prefer the truthful behaviors for the payoff-dominant outcome.
A potential criticism to the above partition-based approach is that we have to give up eliciting ranking information across blocks. One may naturally wonder whether the Isotonic Mechanism can be carefully tailored to elicit additional comparison information among items across partitioned blocks as such information can help address the efficiency of isotonic regression — so long as it can be truthfully elicited. Unfortunately, our second main result gives a negative answer.
@@
Within a significantly more general class of isotonic-regression-based mechanisms (Mechanism 3), any truthful calibration mechanism has to be partition-based (for some choice of item partitions), thus cannot use any ordering information about items across the partitioned blocks.
This result is formally presented in Section 4.1. It illustrates the fundamental trade-off between the statistical efficiency and incentive requirements in designing statistical estimators using data elicited from strategic sources. Hence the only room left for us to optimize the statistical efficiency of our designed mechanism, while retaining its truthfulness, is to optimize the choice of the item partition; this leads to our third main result. As is conceivable, the choice of partition affects the statistical performance of our mechanism. For example, a partition for the instance in Figure 2 can be simply $\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$, which share a single common owner $\{a_{3}\}$, or can be $\{p_{1},p_{2},p_{3}\},\{p_{4}\}$, and $\{p_{5},p_{6}\}$, which share $\{a_{1},a_{2},a_{3}\},\{a_{3}\}$, and $\{a_{3},a_{4}\}$, respectively. To find high-quality owner-item ownership partitions, we define a natural class of criteria to assess the quality of a partition. Within this class of criteria, we obtain the following result on this proposed mechanism:
Main Result 3.
A simple nearly linear-time greedy algorithm can obtain a partition of the owner-item ownership structure for the proposed mechanism that is near-optimal simultaneously for every criterion in the class.
This result is formally stated in Section 4.3. We also show that it is NP-hard to find the exactly optimal partition even under some simple criterion. In contrast, our greedy algorithm is computationally efficient while achieving constant-ratio approximation simultaneously for all the criteria in our considered class. Such simultaneous approximation is a surprisingly nice property of our problem and is a rare phenomenon generally.  [^2] We are aware of only one other situation in approximated majorization for minimizing symmetric convex objectives [13], where there is a simultaneous approximation guarantee but their optimization problem is very different from ours and their ratio is logarithmic in general. This powerful result has recently been used by Banerjee et al. [4] for the welfare guarantee in information design.
Finally, to validate the empirical performance of our mechanism, we applied it to the conference peer review application and conducted a series of experiments on both synthetic data and ICLR review data. Our experiments exhibit that the proposed mechanism enhances estimation performance relative to the original review scores, a benefit observed consistently across various experimental settings.
2 Problem Formulation
This paper considers the review score calibration problem under a network of overlapping ownership. [^3] Besides the motivating example of academic conference review, similar problems with overlapping ownership also exist in reviews for company projects, songs, movies, etc. Hence, our technical sections study an abstract model with generic terms “owner/item”. There are $m$ owners and $n$ items. The $j$-th owner owns a set of items $\mathcal{I}^{j}\subseteq[n]=\{1,2,\dots,n\}$ and let $|\mathcal{I}^{j}|=n_{j}$ [^4] The ownership relation forms a bipartite graph (see e.g., Figure 2), and we formalize their connections in our proofs in Appendix E. For each item $i\in[n]$, let $R_{i}$ be its ground-truth score and
@@ 2 Problem Formulation
be the raw review score given by its reviewers with noise $z_{i}$. As a friendly reminder of our notation, $i$ and $\mathcal{I}$ are used to denote item index and item set, respectively, whereas $j$ is used to index the owner. Denote by $\bm{z}=(z_{1},z_{2},\dots,z_{n}),\bm{R}=(R_{1},R_{2},\dots,R_{n})$ and $\bm{y}=(y_{1},y_{2},\dots,y_{n})$, respectively, the vector of noise, the ground-truth scores, and the raw review scores.
In applications such as conference paper reviews, the ground-truth scores $\bm{R}$ can be interpreted as the mean evaluation scores of the papers perceived across all experts in the community. This item evaluation score is mathematically well defined and always exists, [^5] We remark that paper evaluation score should be carefully distinguished from “true merit” of the paper, which is a vague concept and whose existence might be arguable. Unlike merits, a paper’s true evaluation score is well-defined, and is what peer review procedure as well as statistical methods like ours is trying to estimate. but it is very difficult to be accurately observed during the review phase for various reasons including insufficient resources to have enough reviews and insufficient time to truly observe the items’ impact. This is essentially the motivation of peer review, during which this ground-truth score will be estimated by a few experts sampled from the community. The averaged review scores of these experts form our modeling of the raw review scores $\bm{y}$ — a noisy estimation of ground-truth scores $\bm{R}$. Therefore, the high-level objective of a calibration mechanism $\mathcal{M}$ is to output the adjusted review scores $\widehat{\bm{R}}$ that provide a more accurate estimation of the ground truth scores $\bm{R}$. An advantage of such a calibration mechanism for estimating the community’s expected evaluation of papers is that it is naturally compatible with existing peer review systems.
To tackle the review score calibration problem with the limited supply of qualified reviewers, we resort to the design of owner-assisted calibration mechanisms that elicit and utilize self-evaluation data from the item owners. We abstract this mechanism design problem into the construction of an estimator $\widehat{\mathrm{R}}_{\mathcal{M}}$ for $\bm{R}$ that combines the raw scores $\bm{y}$ with the information elicited from owners $\{\pi^{j}\}_{j=1}^{m}$ to improve the estimates of $\bm{R}$. In this paper, we focus on eliciting each owner’s ranking of their items, where $\pi^{j}$ is a permutation of $1,2,\ldots,n_{j}$ that specifies an ordering of the $j$-th owner’s items in $\mathcal{I}^{j}$. The benefits of this design choice and the possible relaxations are discussed in Section 6. The owner-assisted calibration mechanism then proceeds in the following stages.
  1. 1.
    Before Review: Each owner $j$ observes some information of her items $\mathcal{I}^{j}$, is notified of the calibration mechanism and then asked to report a ranking $\pi^{j}$ of her items $\mathcal{I}^{j}$ [^6] If an owner does not report a ranking, a uniformly random ranking will be used. As will be shown later, this design ensures that rational owners participate in the mechanism. It is also without loss of generality to assume that owners report the full ranking of their items, and we defer the discussion to Remark 2.
  2. 2.
    After Review: Once the raw review score $\bm{y}$ is realized, the mechanism accordingly estimates and outputs the adjusted review scores as $\widehat{\bm{R}}=\widehat{\mathrm{R}}_{\mathcal{M}}(\{\pi^{j}\}_{j=1}^{m};\bm{y})$.
The key challenge of this design problem is to balance the tension between the statistical efficiency and incentive compatibility. To formalize this problem, we adopt the mechanism design approach in modeling the strategic incentives of item owners on the calibration results. That is, we assume each owner has a utility function on their items’ adjusted review scores. Let owner $j$’s utility function be $\mathrm{U}^{j}:\mathbb{R}^{n_{j}}\to\mathbb{R}$ such that each owner $j$ derives utility $\mathrm{U}^{j}\left([\widehat{R}_{i}]_{i\in\mathcal{I}^{j}}\right)$ from adjusted review scores $\widehat{\bm{R}}$, output by the calibration mechanism. We discuss how this utility function based on review scores can be translated to the one based on the final outcomes in Section 6. We employ the Bayesian game framework [15] to study the owners’ strategic decision-making at the information elicitation stage during submission, where the realization of the raw review scores $\bm{y}$ is uncertain to the owners (e.g., due to the unknown reviewer assignments and noise). We say a profile of owners’ report $\{\pi^{j}\}_{j=1}^{m}$ forms a Bayes-Nash Equilibrium (BNE)  [^7] It suffices to consider pure-strategy NEs in most part of this paper and we defer the additional notations for the more general definition of mixed-strategy NEs to Appendix 1. under mechanism $\mathcal{M}$ if for any owner $j\in[m]$, given others’ report $\pi^{-j}=\{\pi^{j^{\prime}}\}_{j^{\prime}\neq j}$, $j$’s expected utility by reporting $\pi^{j}$ is no worse than reporting any other possible ranking $\widetilde{\pi}^{j}$ or, mathematically,
$\quad\mathop{\mathbf{E}}_{\bm{y}}\left[\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\pi^{j},\pi^{-j};\bm{y})\right)\right]\geq\mathop{\mathbf{E}}_{\bm{y}}\left[\mathrm{U}^{j}\left(\widehat{\mathrm{R}}_{\mathcal{M}}(\widetilde{\pi}^{j},\pi^{-j};\bm{y})\right)\right],\quad\forall j\in[m].$ (Equilibrium Condition)
We say a mechanism $\mathcal{M}$ is truthful if every owner $j$ reporting the true ranking $\pi^{j}$ of her items forms a Bayes-Nash equilibrium under $\mathcal{M}$. Through a standard revelation principle argument [29], it is without loss of generality to restrict the design space to truthful mechanisms. This equilibrium condition is sufficient to guarantee the truthfulness against unilateral deviations by each individual owner. We discuss potential extensions to address group manipulation in Section 6.
To demonstrate the theoretical guarantees of our mechanism, we make the following assumptions that are natural in peer review applications. In Section 5, we also conduct empirical studies that demonstrates the robust performance of our mechanism for settings even beyond these assumptions.
Assumption 2.1 (Informed Owners).
For each $j$, the $j$-th owner knows the relative ranking of her own items in $\mathcal{I}^{j}$, according to their ground-truth scores.
Assumption 2.2 (Exchangeable Noise Distribution).
@@
The review noise vector $\bm{z}=(z_{1},\dots,z_{n})$ follows an exchangeable distribution in the sense that $(z_{1},\dots,z_{n})$ has the same probability distribution in $\mathbb{R}^{n}$ as in $\rho\circ\bm{z}:=(z_{\rho(1)},\dots,z_{\rho(n)})$ for any permutation $\rho$ of $1,2,\dots,n$.
Assumption 2.3 (Convex Utility).
@@
Assumption 2.1 is standard in mechanism design and reflects each owner’s private knowledge about their items without requiring any knowledge about other items. [^8] For instance, in basic models of auction design, bidders are assumed to perfectly know their value about the item [29]. Our assumption of owners knowing the ranking information is also inspired by many machine learning applications that learn parameters from humans’ ranking/comparison data which are posited to be more accurate than their knowledge of absolute values [44, 3]. Assumption 2.2 imposes symmetry on review noises, and Assumption 2.3 captures a utility structure that naturally arises in “high-risk-high-reward” tournament settings (Section 6 discusses more about the root of this assumption, its adoption in similar economic problems, and empirical evidences from ICLR data).
3 An Isotonic Mechanism for Completely Overlapping Ownership
We begin with a rank-calibrated score estimator $\widehat{\mathrm{R}}(\pi;\bm{y})$ that employs isotonic regression on raw review scores $\bm{y}$ and any reported ranking $\pi$ from the owner, via the following convex program,